OpenStudy (toxicsugar22):

(x-4)^2/9-(y+2)^2/16=1 give the exact coordinates of any four points on the hyperbola. and show me how you get this

OpenStudy (toxicsugar22):

hi can u help me

OpenStudy (watchmath):

For \(y=-2\) we have \((x-4)^2=9\). Hence\(x-4=\pm 3\). So \(x=4\pm 3\). Then \((7,-2),(1,-2)\)\) are on the hyperbola. For \(y=2\) we have \((x-4)^2/9=2\). Then \((x-4)^2=18\). Hence \(x=4\pm 3\sqrt{2}\). So another two points are \((4+3\sqrt{2},2)\) and \((4-3\sqrt{2},2)\).

OpenStudy (toxicsugar22):

thnaks u can also find these points on the graph

OpenStudy (toxicsugar22):

can u help me with 2 more

OpenStudy (toxicsugar22):

can u help me